Suppose a shipment of 140 electronic components contains 3 defective components. To determine whether the shipment should be​ accepted, a​ quality-control engineer randomly selects 3 of the components and tests them. If 1 or more of the components is​ defective, the shipment is rejected. What is the probability that the shipment is​ rejected? can you help me solve this with statcrunch
Answer by Edwin McCravy(20086)

First we find the probability that it will be accepted and then 
subtract from 1.

There are 3 defective ones, and therefore 140-3 = 137 good ones.

That means the first one selected will be good with a probability of 137/140, and
the second one selected will be good with a probability of 136/139, and
the the third one selected will be good with a probability of 135/138.

Probability that shipment will be accepted = (137/140)(136/139)(135/138) =

0.9355370258

Probability that shipment will be rejected = 1 - 0.9355370258 = 

0.0633629742

Edwin